# Indirectly Parameterized Concrete Autoencoders

### v1

@@ 4.1 Improved Training and Generalization with IP
Dataset
Recon. error
Accuracy
MNIST
3.00$\times$
4.00$\times$
MNIST-Fashion
3.38$\times$
4.60$\times$
ISOLET
3.83$\times$
18.77$\times$
COIL-20
4.15$\times$
25.68$\times$
Smartphone HAR
4.15$\times$
3.70$\times$
Mice Protein
2.53$\times$
12.92$\times$
@@ 4.2 Special Cases of IP-CAE
As addressed in Section 2.4, CAE is a special case of IP-CAE with $\boldsymbol{W}=\mathbf{I}$ and $P=D$. Two additional special cases of IP-CAE that preserve the learning rate scaling properties (Appendix C) in the update rule while offering even simpler formulations are as follows:
@@ 4.3 Hidden Layers
It is worth noting that the main promise of neural network-based embedded feature selection is that it can be non-linear. We observe that for the original CAE, training is unstable with spikes in validation error, hindering a smooth convergence to the optimal solution. The problem worsens for decoders with multiple hidden layers. This is illustrated in Figure 5(b), where we investigate this issue for a linear decoder, an MLP decoder with one hidden layer, and an MLP decoder with two hidden layers. Our improved parameterization results in a smooth descent into a final validation loss that is lower than for the original CAE regardless of decoder complexity, and thus allows for the joint optimization of feature selection and non-linear decoders.
(a) Reconstruction
(b) Classification
@@ 4.4 Comparison with STG and LassoNet
We demonstrate that CAE falls behind STG and LassoNet by a significant margin both for reconstruction and classification. But IP-CAE significantly outperforms them in test accuracy (Table 2) and reconstruction error (Table 1) on all datasets. Additionally, being a stochastic method, CAE is prone to high variance. This problem is reduced in IP-CAE on most datasets, as shown in the figures.
@@ 5 Discussion
In this paper, we addressed the practical challenges of training CAEs. We proposed IP-CAEs which implicitly alleviate redundant features and instability. Our approach achieves state-of-the-art reconstruction error and accuracy for all datasets considered, up to 20x faster than vanilla CAE.
@@
With the goal of understanding IP-CAE’s success, we introduced GJSD regularization which forces unique selections. This baseline significantly improves CAE in every dataset and task, but falls behind IP, as well as LassoNet and STG. This, interestingly, indicates that the effect of IP cannot be solely attributed to removing duplicate features. IP and GJSD are not exclusive and can be combined, but we found adding GJSD regularization to IP-CAE not to improve results significantly.
@@ Acknowledgements
This work was partially supported by KTH Digital Futures, the Swedish Foundation for Strategic Research grants BD15-0043 and ID19-0052, the Swedish eScience Research Centre (SeRC), the Marie Skłodowska-Curie Actions project ”MODELAIR” through grant no. 101072559, and Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The computations were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre as well as Alvis cluster access provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers University of Technology partially funded by the Swedish Research Council through grant agreement no. 2022-06725
@@ Appendix A Experimental Details
In this section, we provide a detailed description of the experimental setup.
@@ A.4 Data and preprocessing
For COIL-20 we use the version of the dataset provided by (Li et al., 2017). For MNIST and Fashion-MNIST, we use the versions provided in Torchvision. For the other datasets, ISOLET, Smartphone HAR, and Mice protein we use the version provided at UCI (Cole & Fanty, 1994; Anguita et al., 2013; Higuera et al., 2015a).
@@
We instead use the imputation method described by the authors of the Mice Protein dataset (Higuera et al., 2015b), which means averaging missing protein expression values with the average expression corresponding to that protein for the same class of mice.
@@ B.6 Runtime
Dataset Classification Reconstruction
@@
Appendix C Update rules for IP-CAE

### v2

@@ 4.1 Improved Training and Generalization with IP
Dataset Recon. error Accuracy
MNIST 3.00$\times$ 4.00$\times$
MNIST-Fashion 3.38$\times$ 4.60$\times$
ISOLET 3.83$\times$ 18.77$\times$
COIL-20 4.15$\times$ 25.68$\times$
Smartphone HAR 4.15$\times$ 3.70$\times$
Mice Protein 2.53$\times$ 12.92$\times$
@@
(a) Reconstruction
(b) Classification
@@ 4.2 Special Cases of IP-CAE
As addressed in Section 2.4, CAE is a special case of IP-CAE with $\boldsymbol{W}=\mathbf{I}$ and $P=D$. Two additional special cases of IP-CAE that preserve the learning rate scaling properties (Appendix D) in the update rule while offering even simpler formulations are as follows:
@@ 4.3 Hidden Layers
It is worth noting that the main promise of neural network-based embedded feature selection is that it can be non-linear. We observe that for the original CAE, training is unstable with spikes in validation error, hindering a smooth convergence to the optimal solution. The problem worsens for decoders with multiple hidden layers. This is illustrated in Figure 6, where we investigate this issue for a linear decoder, an MLP decoder with one hidden layer, and an MLP decoder with two hidden layers. Our improved parameterization results in a smooth descent into a final validation loss that is lower than for the original CAE regardless of decoder complexity, and thus allows for the joint optimization of feature selection and non-linear decoders.
@@ 4.4 Comparison with STG and LassoNet
We demonstrate that CAE falls behind STG and LassoNet by a significant margin both for reconstruction and classification. But IP-CAE significantly outperforms them in test accuracy (Table 2) and reconstruction error (Table 1) on all datasets. Additionally, being a stochastic method, CAE is prone to high variance. This problem is reduced slightly in IP-CAE on most datasets, as shown in Figure 5.
@@ 5 Discussion
In this paper, we addressed the practical challenges of training CAEs. We proposed IP-CAEs which implicitly alleviate redundant features and instability. Our approach achieves state-of-the-art reconstruction error and accuracy for all datasets considered, up to 20 times faster than vanilla CAE.
@@
With the goal of understanding IP-CAE’s success, we introduced GJSD regularization, which forces unique selections. This baseline significantly improves CAE in every dataset and task, but falls behind IP, as well as LassoNet and STG. This, interestingly, indicates that the effect of IP cannot be solely attributed to removing duplicate features. IP and GJSD are not exclusive and can be combined, but we found that adding GJSD regularization to IP-CAE not to improve results significantly.
@@ Acknowledgements
We thank the anonymous reviewers for their valuable feedback and suggestions. This work was partially supported by KTH Digital Futures, the Swedish eScience Research Centre (SeRC), the Swedish Foundation for Strategic Research grants BD15-0043 and ID19-0052, the Marie Skłodowska-Curie Actions project ”MODELAIR” through grant no. 101072559, and Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The computations were enabled by Alvis cluster access provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers University of Technology partially funded by the Swedish Research Council through grant agreement no. 2022-06725 as well as the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre.
@@ Appendix A Experimental Details
In this appendix, we provide a detailed description of the experimental setup.
@@ A.4 Data and preprocessing
For COIL-20 we use the version of the dataset provided by (Li et al., 2017). For MNIST and Fashion-MNIST, we use the versions provided in Torchvision. For the other datasets, ISOLET, Smartphone HAR, and Mice protein we use the version provided at UCI (Fanty & Cole, 1990; Anguita et al., 2013; Higuera et al., 2015).
@@
We instead use the imputation method described by the authors of the Mice Protein dataset (Higuera et al., 2015), which means averaging missing protein expression values with the average expression corresponding to that protein for the same class of mice.
@@ B.6 Runtime
Dataset Classification Reconstr.
@@
Appendix C IP-CAE Updates
In this appendix, we include additional plots showing how components of the IP-CAE update change over time (Figure 10).
(a) $\boldsymbol{\alpha}$
(b) $\psi$
(c) $\boldsymbol{W}$
Figure 10: IP-CAE Update. Components of the update rule related to (a) $\boldsymbol{\alpha}$, (b) $\psi$, and (c) $\boldsymbol{W}$ for each step throughout training for classification on ISOLET. In (a), the value is compared to its equivalent in the vanilla CAE.
Appendix D Update Rules for IP-CAE
In this appendix, we derive the update rules for IP-CAE.